ADNAN MUHAMMED

Grade 12,

Find the period of the function f(x) =[6x+7] +cos(pix)-6x where [.] denotes the greatest integer function.

Find the period of the function f(x) =[6x+7] +cos(pix)-6x where [.] denotes the greatest integer function.

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
[6x+7] = (6x+7) – {6x+7}
        Hence, f(x) = 6x+7 – {6x+7} + cos(pix) – 6x
                            = 7-{6x+7}+cos(pix)
Period of {6x+7} = 1/6
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Period of cos(pix) = 2
Period of 7 = indefinite (constant function)
Therefore , period of f(x) = LCM of 2 and 1/6
                                         = 2.